On an algebra test, the highest grade was 42 points higher than the lowest grade. The sum of two grades was 138. Find the lowest grade
Answer by arKed(27)

N1 + N2 = 138
N2 = 42 + N1
42 + N1 + N1 = 138
42 + 2N1 = 138
2N1 / 2 = (138 -42) / 2
N1 = 48..ANS.
N2 = 42 + N1
N2 = 42 + 48
N2 = 90..ANS.
N1 + N2 = 138
48 + 90 = 138
138 + 138